A customer wants to make a teepee in his backyard for his children. He plans to use lengths of PVC plumbing pip for the supports on the teepee, and he wants the teepee to be 12 feet across and 8 feet tall. How long should the pieces of PVC plumbing pipe be?
Answer by checkley79(3341)

You have a right triangle.
a^2+b^2=c^2
6^2+8^2=c^2
36+84=c^2
120=c^2
c=sqrt120
c=10.95 ft. is the length of the poles (hypotenuse).